OpenStudy (anonymous):

Solve: 3x/(2) + 1/x=3x/4

OpenStudy (bahrom7893):

Okay:\[(3x)/2 + (1/x) = (3x)/4\]

OpenStudy (bahrom7893):

multiply everything by x: \[(3x^2)/2+1=(3x^2)/4\]

OpenStudy (bahrom7893):

Multiply everything by 4: \[6x^2 + 4 = 3x^2\]

OpenStudy (bahrom7893):

Move 3x^2 to the left and 4 to the right: \[6x^2 - 3x^2 = -4\]

OpenStudy (bahrom7893):

\[3x^2 = -4\]

OpenStudy (bahrom7893):

\[x^2 = -4/3\] The solution doesn't exist or just use i

OpenStudy (anonymous):

would it be x= i radical 4/3